OpenStudy (anonymous):

Find all solutions of the equation and express them in the form a + bi. x2 − 10x + 26 = 0 x=________ (positive imaginary) y=________ (negative imaginary)

OpenStudy (anonymous):

apply the quadratic formula

myininaya (myininaya):

y ?

OpenStudy (anonymous):

(x-5)^2 +1 =0

myininaya (myininaya):

oh

myininaya (myininaya):

your esp is better than mine

myininaya (myininaya):

lol

OpenStudy (anonymous):

x=5+-i

OpenStudy (anonymous):

x=5+i , x=5-i

OpenStudy (anonymous):

wait , didnt read lol

OpenStudy (anonymous):

one is for +1 and one is for -1

myininaya (myininaya):

\[x^2-10x=-26\] \[x^2-10x+25=+25-26\] \[(x-5)^2=-1\]\[(x-5)=\pm i\] \[x=5 \pm i\]

OpenStudy (anonymous):

still seems like a bit of a stupid question

OpenStudy (anonymous):

"positive" and "negative" imaginary

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

i dont understand what they are asking really

OpenStudy (anonymous):

x=1, y=-1 its pretty stupid

OpenStudy (anonymous):

i typed it wrong i just realized its x and x

OpenStudy (anonymous):

and its asking for the positive and negative imaginary part

OpenStudy (anonymous):

does that make sense